OpenStudy (anonymous):

let x be the seats in the small van and y the seats in the large van. how many seats does the large van have? a company uses two vans to transport worker from a free parking lot to the workplace between 7:00 and 9:00a.m. one van has 6 more seats than the other. the smaller van makes two trips every morning while the larger one makes only one trip. the two vans can transport 57 people, maximum.

OpenStudy (anonymous):

do u want direct answer or solution?

OpenStudy (anonymous):

direct answer please!

OpenStudy (anonymous):

23

OpenStudy (anonymous):

u sur?

OpenStudy (anonymous):

how come 23?

OpenStudy (anonymous):

dead sure

OpenStudy (anonymous):

Given: 2 vans Given: 1 van #seats = x Given: other van # seats = x + 6 = y Total # trips larger = 1 trip total # trips small = 2 trips max # people = 57 each 2x + y = 57 2x + (x + 6) = 57 3x = 51 x = 51/3 = 17 people y = x + 6 = 17 + 6 = 23 people check: 2x + y = 57 2(17) + 23 = 34 + 23 = 57 ------------------------->OK Therefore: the small van can carry 17 passengers the large van can carry 23 passengers

OpenStudy (anonymous):

let no. of seats in smaller van=x no.of seats in larger=6+x no.of people transported by small in one trip=x in two trips=2x people trans by larger =6+x total ppl tanstd=2x+6x=57

OpenStudy (anonymous):

@hashsam1 has given the correct solution

OpenStudy (anonymous):

yes thanks